# Near Optimal Locality Sensitive Orderings in Euclidean Space

### v2

@@ 4.2 Bichromatic closest pair
Given two sets of points, $R$ and $B$, both subsets of $[0,1)^{d}$, and a parameter $\varepsilon\in(0,1/2)$, one can maintain a $(1+\varepsilon)$-approximation to the closest bichromatic pair $R\times B$. Each insertion/deletion takes $O(\mathcal{E}^{d-1}\log^{2}\mathcal{E}\cdot\log{n})$ time per operation, where $n$ is the maximum size of $|R|+|B|$ over time, and $\mathcal{E}=1/\varepsilon$. This data-structure use $O(\Psi n)$ space, where $\Psi=O(\mathcal{E}^{d-1}\log\mathcal{E})$. At all times, it maintains a pair of points $r\in R$ and $b\in B$, such that $\left\|{rb}\right\|\leq(1+\varepsilon)\mathsf{d}\left({R,B}\right)$, where $\mathsf{d}\left({R,B}\right)=\min_{b\in B,r\in R}\left\|{br}\right\|$.
@@
$\forall p\in R\cup B:r\prec_{\sigma}p\prec_{\sigma}\qquad\implies\qquad\mathsf{d}\left({p,rb}\right)=\min_{q\in rb}\left\|{qp}\right\|\leq(\varepsilon/4)\left\|{rb}\right\|.$
This implies that any pair of points of $P$ in the (close) interval $\sigma[r,b]$, see Eq. (1.1), must be in distance at most $(1+\varepsilon/2)\left\|{rb}\right\|$, from each other, as this is the diameter of the hippodrome $\mathcalb{h}\left({rb,\varepsilon\left\|{rb}\right\|}\right)$. So consider the sorted list of $P\cap\sigma[r,b]$ (according to $\sigma$), and observe that it starts in a point of $R$ and ends in a point of $B$. As such, there must be a consecutive bichromatic pair, and its this appear as an edge in the locality graph. This candidate pair already provides the desired approximation, and the data-structure can only (potentially) return a better one.
@@ 4.3 Dynamic spanners
For a set of points $P$, we show next, that the locality graph $\mathcal{L}=\mathcal{L}(P,\varepsilon/32,1/8)$ is an $(1+\varepsilon)$-spanner. Formally, for two points $x,y\in P$, such that $xy$ is an edge of $\mathcal{L}$, its weight is $\left\|{xy}\right\|$. For any two points $x,y\in P$, let $\mathcalb{g}\left({x,y}\right)$ denote the length of their shortest path in $\mathcal{L}$. The graph $\mathcal{L}$ is $(1+\varepsilon)$-spanner, if for all $x,y\in P$, we have $\mathcalb{g}\left({x,y}\right)\leq(1+\varepsilon)\left\|{xy}\right\|$. In the following, for $X,Y\subseteq P$, let $\mathcalb{g}\left({X,Y}\right)=\min_{x\in X,y\in Y}\mathcalb{g}\left({x,y}\right)$.
@@
Fix a pair $p,q\in P$ and assume by the induction hypothesis that for all pairs $x,y\in P$, such that $\left\|{xy}\right\|<\left\|{pq}\right\|$, we have that $\mathcalb{g}\left({x,y}\right)\leq(1+\varepsilon)\left\|{xy}\right\|$. Let $\sigma\in\Pi$ be the order in $\Pi$ that is $(\varepsilon/32,1/8)$-local for $p$ and $q$,
Figure 4.1:
Let $L=\left\|{pq}\right\|$. The set $P^{\prime}=P\cap\sigma[p,q]$ is contained in the union of the two sets
[…]
where $\gamma=1/8$ and $\xi=\varepsilon/32$, see Figure 4.1 and Figure 1.1 (iii). We consider $P^{\prime}$ to be sorted by $\sigma$. There must be two consecutive points $u,v\in P^{\prime}$, such that $u\in C$ and $v\in D$. By construction, $uv$ is an edge of $\mathcal{L}$. By the gap property, we have that $\left\|{uv}\right\|\geq\mathsf{d}\left({C,D}\right)=(1-2\gamma)\left\|{pq}\right\|\geq(3/4)L$.
Let $u^{\prime},v^{\prime}$ be the projections of $u,v$, respectively, on the line spanned by $pq$. Observe that $\left\|{uu^{\prime}}\right\|\leq\xi L$, and $\left\|{vv^{\prime}}\right\|\leq\xi L$. As such, we have
$\mathcalb{g}\left({p,q}\right)$ $\leq\mathcalb{g}\left({p,u}\right)+\mathcalb{g}\left({u,v}\right)+\mathcalb{g}\left({v,q}\right)\leq(1+\varepsilon)\left\|{pu}\right\|+\left\|{uv}\right\|+(1+\varepsilon)\left\|{vq}\right\|$
$=(1+\varepsilon)(\left\|{pu}\right\|+\left\|{uv}\right\|+\left\|{vq}\right\|)-\varepsilon\left\|{uv}\right\|$
$\leq(1+\varepsilon)(\left\|{pu^{\prime}}\right\|+\left\|{u^{\prime}v^{\prime}}\right\|+\left\|{v^{\prime}q}\right\|+4\xi L)-(3/4)\varepsilon L.$
Observe that $p,u^{\prime},v^{\prime},q$ are all on the same line (spanned by $p$ and $q$), and are contained in the hippodrome $\mathcalb{h}\left({pq,\xi L}\right)$, the point $u^{\prime}$ is closer to $p$ than to $q$, and similarly $v^{\prime}$ is closer to $q$ than to $p$. It follows hat $\left\|{pu^{\prime}}\right\|+\left\|{u^{\prime}v^{\prime}}\right\|+\left\|{v^{\prime}q}\right\|\leq\left\|{pq}\right\|+4\xi L$. Thus, resuming the above calculation, we have
$\mathcalb{g}\left({p,q}\right)$ $\leq(1+\varepsilon)(\left\|{pq}\right\|+8\xi L)-(3/4)\varepsilon L=(1+\varepsilon)(L+8\xi L)-(3/4)\varepsilon L$
$\leq(1+\varepsilon)L+16\xi L-(3/4)\varepsilon L$
$=(1+\varepsilon)L+(1/2-3/4)\varepsilon L\leq(1+\varepsilon)L.$
@@ 5.1 Lower bound of number of LSOs with the $\varepsilon$-local property
For any fixed integer $d$, and $\varepsilon\in(0,1)$, there is a set $P$ of $\Theta(\mathcal{E}^{d})$ points in $\mathbb{R}^{d}$, such that any set of $\Pi$ of LSOs of $P$ with the $\varepsilon$-local property, see Definition 1.1, must be of size $\Omega(\mathcal{E}^{d})$, where $\mathcal{E}=1/\varepsilon$.
@@
Let $m=\left\lfloor{\mathcal{E}/d}\right\rfloor$. Consider the integer grid $P=\left\llbracket m\right\rrbracket^{d}$. Consider any pair of points $p,q\in P$, any other point $z\in P$ is in distance at least one from $p$ or $q$. Setting, $\ell=\left\|{pq}\right\|$, and observing that $\varepsilon\ell\leq m\sqrt{d}/d<1$, we have that
@@ 5.3 Lower bound for LSOs that their locality graph is a spanner
Le et al. [LS21] showed that any $\varepsilon$-spanner in $\mathbb{R}^{d}$ of $n$ points has $\Omega(n/\varepsilon^{d-1})$ edges. Each LSO gives rise to $n-1$ edges in the locality graph. As such, one needs at least $\Omega(1/\varepsilon^{d-1})$ LSOs to construct an $\varepsilon$-spanner using the locality graph.

### v3

@@ 4.2 Bichromatic closest pair
Given two sets of points, $R$ and $B$, both subsets of $[0,1)^{d}$, and a parameter $\varepsilon\in(0,1/2)$, one can maintain a $(1+\varepsilon)$-approximation to the closest bichromatic pair $R\times B$. Each insertion/deletion takes $O(\mathcal{E}^{d-1}\log^{2}\mathcal{E}\cdot\log{n})$ time per operation, where $n$ is the maximum size of $|R|+|B|$ over time, and $\mathcal{E}=1/\varepsilon$. This data-structure use $O(\Psi n)$ space, where $\Psi=O(\mathcal{E}^{d-1}\log\mathcal{E})$. At all times, it maintains a pair of points $r\in R$ and $b\in B$, such that $\left\|{r-b}\right\|\leq(1+\varepsilon)\mathsf{d}\left({R,B}\right)$, where $\mathsf{d}\left({R,B}\right)=\min_{b\in B,r\in R}\left\|{b-r}\right\|$.
@@
$\forall p\in R\cup B:r\prec_{\sigma}p\prec_{\sigma}b\qquad\implies\qquad\mathsf{d}\left({p,rb}\right)=\min_{q\in rb}\left\|{q-p}\right\|\leq(\varepsilon/4)\left\|{r-b}\right\|.$
This implies that any pair of points of $P$ in the (close) interval $\sigma[r,b]$, see Eq. (1.1), must be in distance at most $(1+\varepsilon/2)\left\|{r-b}\right\|$ from each other, as this is the diameter of the hippodrome $\mathcalb{h}\left({rb,\varepsilon\left\|{r-b}\right\|}\right)$. So consider the sorted list of $P\cap\sigma[r,b]$ (according to $\sigma$), and observe that it starts in a point of $R$ and ends in a point of $B$. As such, there must be a consecutive bichromatic pair, and its this appear as an edge in the locality graph. This candidate pair already provides the desired approximation, and the data-structure can only (potentially) return a better one.
@@ 4.3 Dynamic spanners
For a set of points $P$, we show next, that the locality graph $\mathcal{L}=\mathcal{L}(P,\varepsilon/32,1/8)$ is an $(1+\varepsilon)$-spanner. Formally, for two points $x,y\in P$, such that $xy$ is an edge of $\mathcal{L}$, its weight is $\left\|{x-y}\right\|$. For any two points $x,y\in P$, let $\mathcalb{g}\left({x,y}\right)$ denote the length of their shortest path in $\mathcal{L}$. The graph $\mathcal{L}$ is $(1+\varepsilon)$-spanner, if for all $x,y\in P$, we have $\mathcalb{g}\left({x,y}\right)\leq(1+\varepsilon)\left\|{x-y}\right\|$. In the following, for $X,Y\subseteq P$, let $\mathcalb{g}\left({X,Y}\right)=\min_{x\in X,y\in Y}\mathcalb{g}\left({x,y}\right)$.
@@
Fix a pair $p,q\in P$ and assume by the induction hypothesis that for all pairs $x,y\in P$, such that $\left\|{x-y}\right\|<\left\|{p-q}\right\|$, we have that $\mathcalb{g}\left({x,y}\right)\leq(1+\varepsilon)\left\|{x-y}\right\|$. Let $\sigma\in\Pi$ be the order in $\Pi$ that is $(\varepsilon/32,1/8)$-local for $p$ and $q$,
Figure 4.1: Illustration for the proof of Theorem 4.4.
Let $L=\left\|{p-q}\right\|$. The set $P^{\prime}=P\cap\sigma[p,q]$ is contained in the union of the two sets
$C=\mathcalb{b}\left({p,\gamma L}\right)\cap\mathcalb{h}\left({pq,\xi L\bigr{.}}\right)\qquad\text{and}\qquad D=\mathcalb{b}\left({q,\gamma L}\right)\cap\mathcalb{h}\left({pq,\xi L\bigr{.}}\right),$
where $\gamma=1/8$ and $\xi=\varepsilon/32$, see Figure 4.1 and Figure 1.2 (iii). We consider $P^{\prime}$ to be sorted by $\sigma$. There must be two consecutive points $u,v\in P^{\prime}$, such that $u\in C$ and $v\in D$. By construction, $uv$ is an edge of $\mathcal{L}$. By the gap property, we have that
$\left\|{u-v}\right\|\geq\mathsf{d}\left({C,D}\right)=(1-2\gamma)\left\|{p-q}\right\|\geq(3/4)L.$ (4.1)
Let $u^{\prime},v^{\prime}$ be the projections of $u,v$, respectively, on the line spanned by $pq$, see Figure 4.1. We have the following:
  1. (A)
    $\left\|{u-u^{\prime}}\right\|\leq\xi L$ and $\left\|{v-v^{\prime}}\right\|\leq\xi L$.
  2. (B)
    $\left\|{p-u}\right\|\leq\left\|{p-u^{\prime}}\right\|+\left\|{u^{\prime}-u}\right\|\leq\left\|{p-u^{\prime}}\right\|+\xi L$.
  3. (C)
    $\left\|{v-q}\right\|\leq\left\|{v^{\prime}-q}\right\|+\xi L$.
  4. (D)
    $\left\|{u-v}\right\|\leq\left\|{u^{\prime}-v^{\prime}}\right\|+2\xi L$.
  5. (E)
    As $p,u^{\prime},v^{\prime},q$ are all on the same line (spanned by $p$ and $q$), and are contained in the hippodrome $\mathcalb{h}\left({pq,\xi L}\right)$, the point $u^{\prime}$ is closer to $p$ than to $q$, and similarly $v^{\prime}$ is closer to $q$ than to $p$. It follows hat $\left\|{p-u^{\prime}}\right\|+\left\|{u^{\prime}-v^{\prime}}\right\|+\left\|{v^{\prime}-q}\right\|\leq\left\|{p-q}\right\|+4\xi L$.
As such, by induction, we have
$\mathcalb{g}\left({p,q}\right)$ $\leq\mathcalb{g}\left({p,u}\right)+\mathcalb{g}\left({u,v}\right)+\mathcalb{g}\left({v,q}\right)\leq(1+\varepsilon)\left\|{p-u}\right\|+\left\|{u-v}\right\|+(1+\varepsilon)\left\|{v-q}\right\|$
$=\underbrace{(1+\varepsilon)\bigl{(}\left\|{p-u}\right\|+\left\|{u-v}\right\|+\left\|{v-q}\right\|\bigr{)}}_{X}-\varepsilon\left\|{u-v}\right\|.$
We have that
$X$ $\leq(1+\varepsilon)(\left\|{p-u^{\prime}}\right\|+\left\|{u^{\prime}-v^{\prime}}\right\|+\left\|{v^{\prime}-q}\right\|+4\xi L)$
$\leq(1+\varepsilon)(\left\|{p-q}\right\|+8\xi L)$
$=(1+\varepsilon)(L+8\xi L)$
$\leq(1+\varepsilon)L+16\xi L,\leq(1+\varepsilon)L+\smash{\frac{1}{2}}\varepsilon L,$
as $\xi=\varepsilon/32$. The gap property, as stated in Eq. (4.1), implies that $\varepsilon\left\|{u-v}\right\|\geq\tfrac{3}{4}\varepsilon L$, and thus
$\mathcalb{g}\left({p,q}\right)\leq X-\varepsilon\left\|{u-v}\right\|\leq(1+\varepsilon)L+\frac{\varepsilon}{2}L-\frac{3}{4}\varepsilon L\leq(1+\varepsilon)L.$
@@ 5.1 Lower bound of number of LSOs with the $\varepsilon$-local property
For any fixed integer $d$, and $\varepsilon\in(0,1)$, there is a set $P$ of $\Theta(\mathcal{E}^{d})$ points in $\mathbb{R}^{d}$, such that any set $\Pi$ of LSOs of $P$ with the $\varepsilon$-local property, see Definition 1.1, must be of size $\Omega(\mathcal{E}^{d})$, where $\mathcal{E}=1/\varepsilon$.
@@
Let $m=\left\lfloor{\mathcal{E}/d}\right\rfloor$. Consider the integer grid $P=\left\llbracket m\right\rrbracket^{d}$. Consider any pair of points $p,q\in P$, any other point $z\in P$ is in distance at least one from $p$ or $q$. Setting, $\ell=\left\|{p-q}\right\|$, and observing that $\varepsilon\ell\leq m\sqrt{d}/d<1$, we have that
@@ 5.3 Lower bound for LSOs that their locality graph is a spanner
Le and Solomon ls-toes-19,ls-toes-19-arxiv showed that any $\varepsilon$-spanner in $\mathbb{R}^{d}$ of $n$ points has $\Omega(n/\varepsilon^{d-1})$ edges. Each LSO gives rise to $n-1$ edges in the locality graph. As such, one needs at least $\Omega(1/\varepsilon^{d-1})$ LSOs to construct an $\varepsilon$-spanner using the locality graph.